19. A compound with the formula CoH140 shows a strong broad IR signal near 3300 cm! The H NMR spectrum is tabulated below.

Need help proposing a structure given the following data. Thank you
Answer #1

Mole allar formula > G4t Go Hito Degree of Unsaturation - C+1- multiplicity = 2n1+1, 1=112, ne No. g pretros y = 2n1 +1,1=112